# On “confirmatory” methodological research in statistics and related fields

### v4

@@ Abstract
Empirical substantive research, such as in the life or social sciences, is commonly categorized into the two modes exploratory and confirmatory, both of which are essential to scientific progress. The former is also referred to as hypothesis-generating or data-contingent research, while the latter is also called hypothesis-testing research. In the context of empirical methodological research in statistics, however, the exploratory–confirmatory distinction has received very little attention so far. Our paper aims to fill this gap. First, we revisit the concept of empirical methodological research through the lens of the exploratory–confirmatory distinction. Secondly, we examine current practice with respect to this distinction through a literature survey including 115 articles from the field of biostatistics. Thirdly, we provide practical recommendations towards more appropriate design, interpretation, and reporting of empirical methodological research in light of this distinction. In particular, we argue that both modes of research are crucial to methodological progress, but that most published studies—even if sometimes disguised as confirmatory—are essentially exploratory in nature. We emphasize that it may be adequate to consider empirical methodological research as a continuum between “pure” exploration and “strict” confirmation, recommend transparently reporting the mode of conducted research within the spectrum between exploratory and confirmatory, and stress the importance of study protocols written before conducting the study, especially in confirmatory methodological research.
@@ 2.1 Preliminary remarks
Before we review the exploratory–confirmatory distinction in substantive research, we want to make two clarifying remarks on the terminology used in the remainder of this paper. First, perhaps stating the obvious, the hypotheses that are either generated or tested in the two modes of research must be scientific hypotheses, i.e. hypotheses about phenomena one intends to evaluate using a scientific process after stating them [152]. The evaluation of these hypotheses can be rooted in different scientific approaches and is not restricted to particular methods. For example, the use of the word “hypothesis-testing” to describe research does not prescribe that a statistical test must be employed, which leads to our following second remark.
During the discussion of the exploratory–confirmatory distinction and the research process in general, terms such as “hypothesis”, “testing”, “prediction”, “variables”, or “theory” naturally come up. For some readers, these may already be associated with certain concepts due to their specific use in statistics (e.g., in the context of null hypothesis testing or prediction modeling). However, in the context of this paper, they are almost exclusively used with a non-technical, field-independent meaning, referring to overarching scientific concepts and ideas, such that they could be used to describe studies in many disciplines. Thus, unless specifically stated otherwise, these and similar scientific terms are always meant in a broader, non-statistical sense in the remainder of the paper. Finally, let us note that we intentionally use the terms exploratory and confirmatory research and not studies. That is because a single study may have both exploratory and confirmatory objectives, analyses, and results.
@@ 2.2 Exploratory and confirmatory research in substantive disciplines applying statistical methods
Empirical substantive research is commonly categorized into the two modes exploratory and confirmatory research. This pair of labels is also commonly referred to as hypothesis-generating vs. hypothesis-testing, data-contingent research vs. data-independent, or postdiction vs. prediction [97, 132, 138]. Collectively, these four pairs of labels broadly convey the difference between the two modes of research. More detailed characterizations of exploratory and confirmatory research have been offered in publications on psychology [116, 157], biology [122, 136], preclinical [99, 100, 125] and clinical research [120], educational research [107], and linguistics [142], among others. Although addressing audiences in different fields, most of the published definitions share a few key aspects, which we will summarize below. This indicates that there is a sort of consensus with respect to the exploratory–confirmatory distinction in the context of empirical substantive research. It should also be noted that, regardless of the field or assigned labels, the two approaches have distinct but equally valuable purposes, complement each other, and are both essential to scientific progress [133, 136, 138, 142, 145, 157].
@@
The second key aspect distinguishing the two modes of research relates to the data that is used. While there are no restrictions on the data that may be used for exploratory research, confirmatory research often requires the collection of new data [122, 136, 145]. At least, the used data should be different from the data that was used to generate the prespecified hypotheses [116]. Additionally, in confirmatory research, the hypotheses that are prespecified directly inform the research design, data collection or choice, and planning of the data analysis which are similarly specified before the start of the study to ensure a valid and rigorous evaluation. Exploratory research, on the other hand, which intends to explore data more freely to identify patterns, cannot and should not be planned in its entirety at the beginning of a project. Therefore, fewer details about the research design will be specified at that time. However, this does not mean that exploratory research, lacking clear hypotheses, is done aimlessly without any direction or planning. For example, [150] defines exploratory research as a “broad-ranging, purposive, systematic, prearranged undertaking designed to maximize the discovery of generalizations leading to description and understanding” (p. 3).
Another way the distinction has been framed is with the “researcher degrees of freedom” (RDFs; [147]), which refer to the available flexibility in a study’s design, data collection, and analysis. As previously stated, in confirmatory research this flexibility is intentionally restricted (i.e. the RDFs are reduced) to provide sound empirical evidence about clear hypotheses. In contrast, in exploratory research, the RDFs are intentionally utilized to maximize the potential for unexpected discoveries [153, 116].
@@ 2.3 HARKing and preregistration
As mentioned in the last subsection, exploratory research usually gives researchers large freedom in their investigation of data as they look for results that could turn into hypotheses. While this is to some degree required for exploration to fulfill its intended purpose, it also invites the use of questionable research practices (QRPs). A QRP that is particularly related to the exploratory–confirmatory distinction is post hoc theorizing or “hypothesizing after results are known” (HARKing; [124]). A classical example of HARKing would be a researcher retroactively fitting a theory or hypothesis to an interesting exploratory research finding and reporting the combination as if they had stated the theory or hypothesis before looking at the data (i.e. as a confirmatory finding). We want to emphasize that HARKing may also happen unintentionally due to cognitive biases (e.g., confirmation bias or hindsight bias). However, whether intentional or not, presenting an exploratory finding as confirmatory leads to overconfidence in the result. The risk that this result is a false research finding and, therefore, cannot be replicated is higher than if the research were truly confirmatory [92, 138].
@@ 3.2 Exploratory and confirmatory methodological research
Method A is more robust than method B in setting X, and vice-versa for setting Y.
@@
When embarking on confirmatory research, a detailed research protocol with the precise study design including the specific hypotheses to be investigated and the experimental plan as well as the analysis plan must be written before conducting the study. This helps to decrease unwanted data dependency and to increase consistency between all research parts as well as internal validity so that the research actually addresses what is of interest. For research based on real datasets, the protocol should also contain sample size considerations (i.e. a justification of the number of considered real datasets) to ensure adequate statistical power for detecting the effects of interest. In the same vein, the number of simulation repetitions per setting in a simulation study should be chosen to ensure sufficient precision of the results.
@@ 3.3 The case of articles presenting new methods
Given that much of empirical methodological research appears to be published in articles presenting a new method (see the results of our survey in Section 4.1), that is, in articles including method creation contents, such studies are given particular consideration here. We make the argument that, in most cases, claims about new methods based on these experiments are the result of exploratory rather than confirmatory research for the following reasons.
Firstly, we argue that, without any previous empirical evidence about the newly proposed method, one probably lacks the information needed to plan a study in a confirmatory manner. Secondly, even if this likely practical limitation would not exist, the fact that the published empirical study is conducted by the creator(s) of a method gives rise to a number of potential issues that weaken the validity of findings and claims about said method. These issues are related to the non-neutrality of the studies due to creators having a preference for or being more familiar with their new method relative to the other methods in their study [89, 90]. Needless to say, such non-neutrality, which may also have only subconscious influence, can lead to bias in every aspect of a study, from planning and design to analysis and reporting.
Thirdly, another potential problem could be described as data leakage between the method creation process and the published empirical study. It occurs when the datasets (or simulation scenarios) analyzed in the empirical study were already utilized in some form during the method creation, for example, in early experiments that guided the creation process, essentially meaning that the same data are used for both method creation and method evaluation [123]. If a dataset used to develop a method is also used to evaluate its performance, the resulting evaluation is likely to be overly optimistic. This is because the method becomes overly tailored to the specific characteristics of that dataset and may not generalize well to others.
Data leakage can also take more subtle and indirect forms. Without involving the datasets directly, leakage may also be present if a method creator had prior knowledge about the datasets (e.g., their characteristics or results of existing studies), which likely biased their method creation, even if it was not consciously utilized [143]. While papers exclusively investigating existing methods (i.e. methods for which an evaluation has been published) can certainly also have neutrality issues, claims therein are potentially less exploratory as information about the methods was already available, and authors often have a weaker personal connection to the methods.
@@ 4.1 A survey of recent Biometrical Journal and Statistics in Medicine articles
We extracted all eligible articles from the issue with which this target was met, resulting in slightly more than 50 articles per journal in our survey. The data on which the results presented in this section are based (i.e. the spreadsheet documenting the survey) is provided in the Supporting Information. Ignoring editorials, comments, tutorials, book reviews, corrections, and letters, we extracted 122 research articles (61 from each journal), of which 115 contain empirical methodological research as defined above ($57$ from BiomJ and $58$ from StatMed). In the vast majority of the surveyed articles ($100/115=87\%$), only simulated data was used to study the investigated method(s). $11$ articles presented real-data methodological studies, and $4$ articles used both simulated data and real data. In articles where the methodological study did not involve real data, numerical results were almost always augmented with an illustration of the method(s) on one or two real datasets.
@@ 4.2 Empirical methodological literature dominated by exploratory research?
As far as real-data studies are concerned, two common practices suggest that they generally do not meet the standards of confirmatory research: a) most studies use only one or a few datasets, and b) due to the limited number of datasets available, most studies at least partially overlap in the datasets they use. [129] surveyed 215 ML papers published between 2008 and 2010 and found that over 80% of studies used either one dataset (25.5%) or 2–10 datasets (55.5%). They also reported that in 63.9% of studies, the datasets were sourced from the popular UCI repository (https://archive.ics.uci.edu/), with some datasets being studied particularly often (e.g., 15 datasets were used in at least ten of the surveyed papers, five of which even in at least 20 papers). Similarly, in their survey of 55 articles with real-data comparison studies (published between 2010 and 2012), [89] reported regarding the number of datasets that the median was 5, the upper quartile was 7.5, and the maximum was 21. More recently, [127] surveyed 103 papers published at major ML conferences between 2016 and 2021 and found that more than 65% of them used three datasets or fewer, while the overall mean number of datasets was 4.1. In the same vein, [126] analyzed the usage of 2,063 datasets across 26,691 ML papers published between 2015 and 2020 and found a concentration on a limited number of datasets within different ML communities that has been increasing in the considered period. Even if one assumes that many studies have prespecified hypotheses, the majority of studies use (very) small numbers of datasets, and the frequent dataset reuse increases the chances of the hypotheses being tested on (partially) the same data that were used to generate them (i.e. data leakage between studies).
It is difficult, perhaps even impossible, to ascertain whether a finding in the existing literature was the result of confirmatory research, but it is probable that the vast majority of the empirical methodological literature so far has presented non-confirmatory research. At this point, we want to reemphasize that exploratory research is just as valuable as confirmatory research, and there are also indications that current empirical methodological research is at least not entirely exploratory. Even though none of the articles included in our survey have explicitly prespecified hypotheses, many of them contain wording that suggests that the authors did have at least vague hypotheses before conducting their study (e.g., frequent use of “as expected” when describing results). However, considering our previous observations about current practices, it is unlikely that findings in those articles would qualify as confirmatory, even if the authors had explicitly prespecified their expectations as hypotheses.
@@ 4.3 Pseudo-confirmatory findings
We denote as pseudo-confirmatory findings non-confirmatory findings that are presented as if they were the result of confirmatory research. Pseudo-confirmatory findings can arise either from research intended as exploratory from the start or from research intended as confirmatory but where QRPs were employed. The term is new, but the phenomenon is known in many fields. Based on our survey, we suspect that pseudo-confirmatory findings are fairly common in empirical methodological research.
@@
In benchmark studies using real datasets, for example, researchers sometimes engage in post hoc exclusions of certain benchmarked methods or datasets. It is also likely that they sometimes exploit the flexibility of the design of their benchmark and perform many different benchmark variations in the hope of finding the superiority of a particular method [90, 134, 137]. Similar mechanisms may be at work in simulation studies [155, 140]. While most authors engaging in QRPs do so unintentionally (i.e. without malicious intention), one may argue that such bad research practices should and could simply be avoided in confirmatory research.
@@ 6 Discussion
In the previous sections, we entertained that in most empirical methodological research exploration and confirmation are not clearly distinguished. In particular, we demonstrated with a survey that in many instances empirical methodological research can be considered what we call pseudo-confirmatory research. This places an epistemic burden on the fields and explains to some extent why parts of computational research are non-replicable and unreliable, and why we need more actual confirmatory methodological research to provide more reliable guidance to applied researchers who want to use these methods. In the previous section we provided recommendations to improve the situation.
@@
It should be noted that the concept of confirmatory research is tied to a set of important and interrelated epistemic questions, the discussion of which is beyond the scope of this paper. Some of these issues have been addressed in a related position paper on empirical research in machine learning coauthored by some of us [114] and other articles. For example, in a discussion for the special collection “Towards Neutral Comparison Studies in Methodological Research” in the Biometrical Journal, [151] point out that asking “Which method is best for a given set of datasets?” may be the wrong question to ask in the first place and “argue that this research question implies assumptions which [they] do not consider warranted in methodological research” (p. 1). Among other things, this is the case because it is very difficult to define the population of datasets the results of a confirmatory study are intended to generalize to. This is a rather pronounced perspective on the important question of how far the results of a benchmark experiment in methodological research can be generalized beyond the specific experimental setup, which is the key goal of confirmatory research.
@@
Finally, especially in psychology and in the meta-scientific literature, some argue that a lack of background theory prevents specifying hypotheses a priori and, hence, that the field is not mature enough for strict confirmatory research [^3] An illustrative example of a strict confirmatory study in this sense, i.e. where the hypothesis being tested is derived from a strong (in this case, physical) theory and basically every experimental detail is prespecified (and preregistered), is the study by [80], who show that “[f]air coins tend to land on the same side they started” (p. 1). [106], which to some extent also holds for machine learning [114]. While it is important to discuss these and related epistemic issues from a meta-scientific perspective, they should not distract from the fact that confirmatory research is possible—at least to an extent that makes it practically useful and valuable—even without a strong theory to derive hypotheses from. In medicine, too, there are sometimes no strong theories to support the hypotheses being tested. Nevertheless, phase III clinical trials are clearly prime examples of confirmatory research. The crucial point is that they are based on successive stages of prior (exploratory) investigations conducted in the different phases of medical research. In this sense, the field of medicine is certainly “mature enough” for confirmatory research even in cases without strong theories. In our opinion, this is not yet fully the case for empirical methodological research, and it should also mature in this sense. With this work, we hope to contribute to this improvement by raising awareness of the conceptual and practical differences between exploration and confirmation in empirical methodological research.

### v5

@@ Abstract
Empirical substantive research, such as in the life or social sciences, is commonly categorized into the two modes exploratory and confirmatory, both of which are essential to scientific progress. The former is also referred to as hypothesis-generating or data-contingent research, while the latter is also called hypothesis-testing research. In the context of empirical methodological research in statistics, however, the exploratory–confirmatory distinction has received very little attention so far. Our paper aims to fill this gap. First, we revisit the concept of empirical methodological research through the lens of the exploratory–confirmatory distinction. Second, we examine current practice with respect to this distinction through a literature survey including 115 articles from the field of biostatistics. Third, we provide practical recommendations toward a more appropriate design, interpretation, and reporting of empirical methodological research in light of this distinction. In particular, we argue that both modes of research are crucial to methodological progress, but that most published studies—even if sometimes disguised as confirmatory—are essentially exploratory in nature. We emphasize that it may be adequate to consider empirical methodological research as a continuum between “pure” exploration and “strict” confirmation, recommend transparently reporting the mode of conducted research within the spectrum between exploratory and confirmatory, and stress the importance of study protocols written before conducting the study, especially in confirmatory methodological research.
@@ 2.1 Preliminary Remarks
Before we review the exploratory–confirmatory distinction in substantive research, we want to make two clarifying remarks on the terminology used in the remainder of this paper. First, perhaps stating the obvious, the hypotheses that are either generated or tested in the two modes of research must be scientific hypotheses, that is, hypotheses about phenomena one intends to evaluate using a scientific process after stating them [152]. The evaluation of these hypotheses can be rooted in different scientific approaches and is not restricted to particular methods. For example, the use of the word “hypothesis-testing” to describe research does not prescribe that a statistical test must be employed, which leads to our following second remark.
During the discussion of the exploratory–confirmatory distinction and the research process in general, terms such as “hypothesis”, “testing”, “prediction”, “variables”, or “theory” naturally come up. For some readers, these may already be associated with certain concepts due to their specific use in statistics (e.g., in the context of null hypothesis testing or prediction modeling). However, in the context of this paper, they are almost exclusively used with a nontechnical, field-independent meaning, referring to overarching scientific concepts and ideas, such that they could be used to describe studies in many disciplines. Thus, unless specifically stated otherwise, these and similar scientific terms are always meant in a broader, nonstatistical sense in the remainder of the paper. Finally, let us note that we intentionally use the terms exploratory and confirmatory research and not studies. That is because a single study may have both exploratory and confirmatory objectives, analyses, and results.
@@ 2.2 Exploratory and Confirmatory Research in Substantive Disciplines Applying Statistical Methods
Empirical substantive research is commonly categorized into the two modes exploratory and confirmatory research. This pair of labels is also commonly referred to as hypothesis-generating versus hypothesis-testing, data-contingent research versus data-independent, or postdiction versus prediction [97, 132, 138]. Collectively, these four pairs of labels broadly convey the difference between the two modes of research. More detailed characterizations of exploratory and confirmatory research have been offered in publications on psychology [116, 157], biology [122, 136], preclinical [99, 100, 125] and clinical research [120], educational research [107], and linguistics [142], among others. Although addressing audiences in different fields, most of the published definitions share a few key aspects, which we will summarize below. This indicates that there is a sort of consensus with respect to the exploratory–confirmatory distinction in the context of empirical substantive research. It should also be noted that, regardless of the field or assigned labels, the two approaches have distinct but equally valuable purposes, complement each other, and are both essential to scientific progress [133, 136, 138, 142, 145, 157].
@@
The second key aspect distinguishing the two modes of research relates to the data that is used. While there are no restrictions on the data that may be used for exploratory research, confirmatory research often requires the collection of new data [122, 136, 145]. At least, the used data should be different from the data that was used to generate the prespecified hypotheses [116]. Additionally, in confirmatory research, the hypotheses that are prespecified directly inform the research design, data collection or choice, and planning of the data analysis, which are similarly specified before the start of the study to ensure a valid and rigorous evaluation. Exploratory research, on the other hand, which intends to explore data more freely to identify patterns, cannot and should not be planned in its entirety at the beginning of a project. Therefore, fewer details about the research design will be specified at that time. However, this does not mean that exploratory research, lacking clear hypotheses, is done aimlessly without any direction or planning. For example, [150] defines exploratory research as a “broad-ranging, purposive, systematic, prearranged undertaking designed to maximize the discovery of generalizations leading to description and understanding” (p. 3).
Another way the distinction has been framed is with the “researcher degrees of freedom” (RDFs; [147]), which refer to the available flexibility in a study’s design, data collection, and analysis. As previously stated, in confirmatory research this flexibility is intentionally restricted (i.e., the RDFs are reduced) to provide sound empirical evidence about clear hypotheses. In contrast, in exploratory research, the RDFs are intentionally utilized to maximize the potential for unexpected discoveries [153, 116].
@@ 2.3 HARKing and Preregistration
As mentioned in the last subsection, exploratory research usually gives researchers large freedom in their investigation of data as they look for results that could turn into hypotheses. While this is to some degree required for exploration to fulfill its intended purpose, it also invites the use of questionable research practices (QRPs). A QRP that is particularly related to the exploratory–confirmatory distinction is post hoc theorizing or “hypothesizing after results are known” (HARKing; [124]). A classical example of HARKing would be a researcher retroactively fitting a theory or hypothesis to an interesting exploratory research finding and reporting the combination as if they had stated the theory or hypothesis before looking at the data (i.e., as a confirmatory finding). We want to emphasize that HARKing may also happen unintentionally due to cognitive biases (e.g., confirmation bias or hindsight bias). However, whether intentional or not, presenting an exploratory finding as confirmatory leads to overconfidence in the result. The risk that this result is a false research finding and, therefore, cannot be replicated is higher than if the research were truly confirmatory [92, 138].
@@ 3.2 Exploratory and Confirmatory Methodological Research
Method A is more robust than method B in setting X, and vice versa for setting Y.
@@
When embarking on confirmatory research, a detailed research protocol with the precise study design, including the specific hypotheses to be investigated and the experimental plan, as well as the analysis plan must be written before conducting the study. This helps to decrease unwanted data dependency and to increase consistency between all research parts as well as internal validity so that the research actually addresses what is of interest. For research based on real datasets, the protocol should also contain sample size considerations (i.e., a justification of the number of considered real datasets) to ensure adequate statistical power for detecting the effects of interest. In the same vein, the number of simulation repetitions per setting in a simulation study should be chosen to ensure sufficient precision of the results.
@@ 3.3 The Case of Articles Presenting New Methods
Given that much of empirical methodological research appears to be published in articles presenting a new method (see the results of our survey in Section 4.1), that is, in articles including method creation contents, such studies are given particular consideration here. We make the argument that, in most cases, claims about new methods based on these experiments are the result of exploratory rather than confirmatory research for the following reasons:
First, we argue that, without any previous empirical evidence about the newly proposed method, one probably lacks the information needed to plan a study in a confirmatory manner. Second, even if this likely practical limitation would not exist, the fact that the published empirical study is conducted by the creator(s) of a method gives rise to a number of potential issues that weaken the validity of findings and claims about said method. These issues are related to the non-neutrality of the studies due to creators having a preference for or being more familiar with their new method relative to the other methods in their study [89, 90]. Needless to say, such non-neutrality, which may also have only subconscious influence, can lead to bias in every aspect of a study, from planning and design to analysis and reporting.
Third, another potential problem could be described as data leakage between the method creation process and the published empirical study. It occurs when the datasets (or simulation scenarios) analyzed in the empirical study were already utilized in some form during the method creation, for example, in early experiments that guided the creation process, essentially meaning that the same data are used for both method creation and method evaluation [123]. If a dataset used to develop a method is also used to evaluate its performance, the resulting evaluation is likely to be overly optimistic. This is because the method becomes overly tailored to the specific characteristics of that dataset and may not generalize well to others.
Data leakage can also take more subtle and indirect forms. Without involving the datasets directly, leakage may also be present if a method creator had prior knowledge about the datasets (e.g., their characteristics or results of existing studies), which likely biased their method creation, even if it was not consciously utilized [143]. While papers exclusively investigating existing methods (i.e., methods for which an evaluation has been published) can certainly also have neutrality issues, claims therein are potentially less exploratory as information about the methods was already available, and authors often have a weaker personal connection to the methods.
@@ 4.1 A Survey of Recent Biometrical Journal and Statistics in Medicine Articles
We extracted all eligible articles from the issue with which this target was met, resulting in slightly more than 50 articles per journal in our survey. The data on which the results presented in this section are based (i.e., the spreadsheet documenting the survey) is provided in the Supporting Information. Ignoring editorials, comments, tutorials, book reviews, corrections, and letters, we extracted 122 research articles (61 from each journal), of which 115 contain empirical methodological research as defined above ($57$ from BiomJ and $58$ from StatMed). In the vast majority of the surveyed articles ($100/115=87\%$), only simulated data was used to study the investigated method(s). Eleven articles presented real-data methodological studies, and $4$ articles used both simulated data and real data. In articles where the methodological study did not involve real data, numerical results were almost always augmented with an illustration of the method(s) on one or two real datasets.
@@ 4.2 Empirical Methodological Literature Dominated by Exploratory Research?
As far as real-data studies are concerned, two common practices suggest that they generally do not meet the standards of confirmatory research: a) most studies use only one or a few datasets, and b) due to the limited number of datasets available, most studies at least partially overlap in the datasets they use. [129] surveyed 215 ML papers published between 2008 and 2010 and found that over 80% of studies used either one dataset (25.5%) or 2–10 datasets (55.5%). They also reported that in 63.9% of studies, the datasets were sourced from the popular UCI repository (https://archive.ics.uci.edu/), with some datasets being studied particularly often (e.g., 15 datasets were used in at least ten of the surveyed papers, five of which even in at least 20 papers). Similarly, in their survey of 55 articles with real-data comparison studies (published between 2010 and 2012), [89] reported regarding the number of datasets that the median was 5, the upper quartile was 7.5, and the maximum was 21. More recently, [127] surveyed 103 papers published at major ML conferences between 2016 and 2021 and found that more than 65% of them used three datasets or fewer, while the overall mean number of datasets was 4.1. In the same vein, [126] analyzed the usage of 2063 datasets across 26,691 ML papers published between 2015 and 2020 and found a concentration on a limited number of datasets within different ML communities that has been increasing in the considered period. Even if one assumes that many studies have prespecified hypotheses, the majority of studies use (very) small numbers of datasets, and the frequent dataset reuse increases the chances of the hypotheses being tested on (partially) the same data that were used to generate them (i.e., data leakage between studies).
It is difficult, perhaps even impossible, to ascertain whether a finding in the existing literature was the result of confirmatory research, but it is probable that the vast majority of the empirical methodological literature so far has presented nonconfirmatory research. At this point, we want to reemphasize that exploratory research is just as valuable as confirmatory research, and there are also indications that current empirical methodological research is at least not entirely exploratory. Even though none of the articles included in our survey have explicitly prespecified hypotheses, many of them contain wording that suggests that the authors did have at least vague hypotheses before conducting their study (e.g., frequent use of “as expected” when describing results). However, considering our previous observations about current practices, it is unlikely that findings in those articles would qualify as confirmatory, even if the authors had explicitly prespecified their expectations as hypotheses.
@@ 4.3 Pseudo-Confirmatory Findings
We denote as pseudo-confirmatory findings nonconfirmatory findings that are presented as if they were the result of confirmatory research. Pseudo-confirmatory findings can arise either from research intended as exploratory from the start or from research intended as confirmatory but where QRPs were employed. The term is new, but the phenomenon is known in many fields. Based on our survey, we suspect that pseudo-confirmatory findings are fairly common in empirical methodological research.
@@
In benchmark studies using real datasets, for example, researchers sometimes engage in post hoc exclusions of certain benchmarked methods or datasets. It is also likely that they sometimes exploit the flexibility of the design of their benchmark and perform many different benchmark variations in the hope of finding the superiority of a particular method [90, 134, 137]. Similar mechanisms may be at work in simulation studies [155, 140]. While most authors engaging in QRPs do so unintentionally (i.e., without malicious intention), one may argue that such bad research practices should and could simply be avoided in confirmatory research.
@@ 6 Discussion
In the previous sections, we entertained that in most empirical methodological research exploration and confirmation are not clearly distinguished. In particular, we demonstrated with a survey that in many instances empirical methodological research can be considered what we call pseudo-confirmatory research. This places an epistemic burden on the fields and explains to some extent why parts of computational research are non-replicable and unreliable, and why we need more actual confirmatory methodological research to provide more reliable guidance to applied researchers who want to use these methods. In the previous section, we provided recommendations to improve the situation.
@@
It should be noted that the concept of confirmatory research is tied to a set of important and interrelated epistemic questions, the discussion of which is beyond the scope of this paper. Some of these issues have been addressed in a related position paper on empirical research in ML coauthored by some of us [114] and other articles. For example, in a discussion for the special collection “Towards Neutral Comparison Studies in Methodological Research” in the Biometrical Journal, [151] point out that asking “Which method is best for a given set of datasets?” may be the wrong question to ask in the first place and “argue that this research question implies assumptions which [they] do not consider warranted in methodological research” (p. 1). Among other things, this is the case because it is very difficult to define the population of datasets the results of a confirmatory study are intended to generalize to. This is a rather pronounced perspective on the important question of how far the results of a benchmark experiment in methodological research can be generalized beyond the specific experimental setup, which is the key goal of confirmatory research.
@@
Finally, especially in psychology and in the meta-scientific literature, some argue that a lack of background theory prevents specifying hypotheses a priori and, hence, that the field is not mature enough for strict confirmatory research [^3] An illustrative example of a strict confirmatory study in this sense, that is, where the hypothesis being tested is derived from a strong (in this case, physical) theory and basically every experimental detail is prespecified (and preregistered), is the study by [80], who show that “[f]air coins tend to land on the same side they started” (p. 1). [106], which to some extent also holds for ML [114]. While it is important to discuss these and related epistemic issues from a meta-scientific perspective, they should not distract from the fact that confirmatory research is possible—at least to an extent that makes it practically useful and valuable—even without a strong theory to derive hypotheses from. In medicine, too, there are sometimes no strong theories to support the hypotheses being tested. Nevertheless, phase III clinical trials are clearly prime examples of confirmatory research. The crucial point is that they are based on successive stages of prior (exploratory) investigations conducted in the different phases of medical research. In this sense, the field of medicine is certainly “mature enough” for confirmatory research even in cases without strong theories. In our opinion, this is not yet fully the case for empirical methodological research, and it should also mature in this sense. With this work, we hope to contribute to this improvement by raising awareness of the conceptual and practical differences between exploration and confirmation in empirical methodological research.